# Massive Digital Over-the-Air Computation for Communication-Efficient Federated Edge Learning

### v1

@@ Abstract
Over-the-air computation (AirComp) is a promising technology converging communication and computation over wireless networks, which can be particularly effective in model training, inference, and more emerging edge intelligence applications. AirComp relies on uncoded transmission of individual signals, which are added naturally over the multiple access channel thanks to the superposition property of the wireless medium. Despite significantly improved communication efficiency, how to accommodate AirComp in the existing and future digital communication networks, that are based on discrete modulation schemes, remains a challenge. This paper proposes a massive digital AirComp (MD-AirComp) scheme, that leverages an unsourced massive access protocol, to enhance compatibility with both current and next-generation wireless networks. MD-AirComp utilizes vector quantization to reduce the uplink communication overhead, and employs shared quantization and modulation codebooks. At the receiver, we propose a near-optimal approximate message passing-based algorithm to compute the model aggregation results from the superposed sequences, which relies on estimating the number of devices transmitting each code sequence, rather than trying to decode the messages of individual transmitters. We apply MD-AirComp to the federated edge learning (FEEL), and show that it significantly accelerates FEEL convergence compared to state-of-the-art while using the same amount of communication resources. To support further research and ensure reproducibility, we have made our code available at https://github.com/liqiao19/MD-AirComp.
@@
Artificial intelligence of things (AIoT), digital over-the-air computation, unsourced random access, federated edge learning, distributed optimization.
@@ I Introduction
With the advancement of the Artificial Intelligence of Things (AIoT), a growing volume of data is being generated by a massive number of Internet-of-Things (IoT) devices, which fuels the ongoing revolution in machine learning (ML) and artificial intelligence (AI) [2, 3]. Traditionally, data collected by IoT devices is offloaded to the cloud or data centers to train massively large models [4]. However, due to concerns over data privacy and the increasing computational capabilities of edge devices like smartphones, AIoT in wireless networks is transitioning away from centralized deep learning towards distributed learning frameworks [5, 4, 7, 6]. One of these emerging frameworks is federated edge learning (FEEL), where a central server coordinates multiple AIoT devices to train an ML model using their respective local datasets and computing resources [8]. As a result, FEEL requires solving the parallel and distributed optimization problem while considering the limited shared wireless resources and interference between devices. The model aggregation process in FEEL involves the repeated uplink (UL) transmission of high-dimensional local gradients or model updates from tens to hundreds of devices, which causes a heavy burden on the multiple access networks [9].
@@ I-A Related Work
To address this problem, the concept of over-the-air computation (AirComp) has been introduced as a means of achieving communication-efficient model aggregation [8, 9, 10]. In the context of model aggregation, the objective is to compute the average of all the local model updates, rather than decoding each transmitted message individually. In analog AirComp, the fundamental principle involves generating and utilizing inter-user interference over the multiple-access channel (MAC). More specifically, every element of the local model update at a device goes through pre-equalization using the inverse of the UL channel gain from the device to the BS acting as the parameter server. Subsequently, the pre-equalized elements are modulated onto the transmit waveform’s amplitude. Through simultaneous transmission from multiple devices, the receiver can directly obtain the sum of the local model updates from multiple devices by leveraging the superposed waveform over the MAC. To further reduce the communication overhead, by exploiting sparsification and error accumulation, the dimension of the local model updates are reduced before analog transmission in [10]. After analog AirComp, the receiver reconstructs the average of local model updates by using the approximate message passing (AMP) algorithm. A higher FEEL accuracy is achieved with limited communication resources in [11], by exploiting time-correlated sparsification on the local model updates. Multi-input multi-output (MIMO)-based AirComp is studied in [12]. Random projection technique proposed in [10] is extended to MIMO channel in [13]. Note that all the participating devices are synchronized in time in [10, 8, 9, 11, 12, 13]. If the synchronization is imperfect, [14] and [15] considered misaligned analog AirComp to solve the problem. Dynamic scheduling of devices is proposed in [16] to optimize the energy efficiency of AirComp in FEEL. We refer the readers to [17] and [18] for a comprehensive overview of FEEL and AirComp. We should emphasize that AirComp is a general technology that turns wireless medium into a joint communication and computation platform. Its benefits can also be applied to other parallel and distributed optimization applications [19]. For instance, AirComp has recently been used for efficient inference at the edge [20, 21].
@@
To fill in this gap, we propose a digital scheme that jointly enables random access and computation among massive AIoT devices. We first revisit grant-free random access (GFRA) protocols, which have been recently studied to enable massive IoT access with low signaling overhead and latency [35]. Since devices can directly transmit their signals without the permission of the base station (BS), GFRA protocols reduce the signaling overhead. There are mainly two types of GFRA schemes, i.e., sourced massive access (SMA) [36, 37, 38, 39, 40] and unsourced massive access (UMA) [42, 41, 44, 43]. Both SMA and UMA schemes adopt non-orthogonal random access codebook for reducing the communication overhead. In SMA, each device has its unique non-orthogonal preamble indicating its identity. The BS performs active device detection and channel estimation, followed by data decoding [37, 38, 39, 40]. In UMA, instead, all the devices adopt the same non-orthogonal codebook. Transmitted bits are modulated by the index of the UMA codewords and decoded at the BS. The goal of the BS is to decode the list of transmitted messages, but not the identities of the active devices [42, 44, 41, 43]. We remark that the identities of active devices are not necessary for model aggregation in FEEL either. Inspired by UMA, our target is to redesign the modulation and decoding modules of conventional UMA for efficient AirComp across a large number of devices.
@@ II System Model
where $\eta$ is the global learning rate. Finally, the updated parameter vector ${\bf w}^{t+1}\in\mathbb{R}^{W}$ is sent back to the devices. Steps (2), (3), and (4) are iterated until a convergence condition is met.
@@
One of the reasons that $K_{a}\ll K$ in FEEL, e.g., in literature [10, 26, 18], is the heterogeneous computation speeds of different devices [48]. For a given time slot, only the devices that already finished their computation task will transmit. Devices’ availability can be another reason, that devices can participate or leave at particular frames, as considered in typical IoT scenarios [2].
@@ III-A1 Quantization Design
We will employ a single quantization codebook to be used by all the devices, denoted by ${\bf U}\in\mathbb{R}^{Q\times N}$, where $N=2^{J}$ denotes $J$-bit quantization with $N$ quantization codewords, $Q\geq 1$ ($Q\in\mathbb{N}$) denotes the length of each quantization codeword, i.e., ${\bf u}_{n}\in\mathbb{R}^{Q}$, $\forall n\in[N]$. Note that $Q=1$ and $Q>1$ indicate scalar quantization and VQ, respectively.
@@ III-A2 Codebook-Based Modulation
We consider the same random access codebook shared by all the devices, denoted by ${\bf P}\in\mathbb{C}^{L\times N}$, where each column of ${\bf P}$ is a codeword of length $L$, and there are $N$ codewords in total. In the proposed scheme, there is a one-to-one mapping between ${\bf P}$ and ${\bf U}$. To be specific, if the device acquires the quantized value $[{\bf U}]_{:,n}$, then the modulation codeword (or sequence) $[{\bf P}]_{:,n}$ that has the same index $n$ will be transmitted. The transmission process is similar to the preamble transmission of random access procedure in 5G communication systems [22]. Hence, the proposed codebook-based modulation is more suitable for the current wireless networks compared to the analog modulation used in AirComp.
@@ III-A3 UL Massive Access Model
We consider a time division duplex system, where the UL and downlink (DL) channels are commonly considered the same due to channel reciprocity [38, 43]. Before each UL communication round of FEEL, the BS first broadcasts a pilot signal and each active device estimates its DL channel based on the pilot signal. Then, the active devices start their grant-free UL transmissions with pre-equalization in a synchronized manner [38]. For the $d$-th element of ${\bf b}_{k}$, $\forall k,d$, the received signal ${\bf Y}_{d}\in\mathbb{C}^{L\times M}$ at the BS can be expressed as
@@
As shown in (7), the noise power can be defined by the SNR and the average power of the received signal, which is commonly used in massive access system models [35]. If the wireless channel $[{\bf h}_{k}^{\rm UL}]_{1}$ is in deep fading, the power of $[{\bf h}_{k}^{\rm UL}]_{1}$ can be extremely small. In this case, pre-equalization might increase the received signal power at the rest of the antennas. As a result, SNR at the first antenna can be too small to conduct proper signal detection. In addition, if the channel power is too small, the transmit power needed for pre-equalization might be too large. Hence, we set a threshold $\epsilon_{h}$ on the channel condition, such that if $|[{\bf h}_{k}]_{1}|<\epsilon_{h}$, the device will not participate in this training round.
@@ IV-B Approximate Message Passing (AMP) Update Rules
where symbol $\approx$” is due to the approximation, the likelihood function and the normalization term are respectively denoted as
@@ IV-C2 Estimation of the activity indicators
For example, to estimate the activity indicator $a_{n,m}$, $\forall n,m$, we partially differentiate $\mathbb{E}\left\{{{\rm ln}~{}p\left({{\bf X,Y}}\right)|{\bf Y}}\right\}$ with respect to $a_{n,m}$ and let it equal zero [55]. According to (IV-B), the posterior distribution $p\left(x_{n,m}|[{\bf Y}]_{:,m}\right)$ can be approximated as $p\left(x_{n,m}|r_{n,m}\right)$, thereby simplifying the complexity of EM estimation. Hence, according to (23) and (24), the sparsity indicators $a_{n,1}$ and $a_{n,m}$, $m\in\{2,3,...,M\}$, can be obtained as
@@ VI-D Results of Wireless Transmission
From Fig. 5, we can observe that the NMSE decreases due to the increase in the sparsity ratio. In addition, the NMSE of “MD-AirComp (6, 20, 15), $M=8$” is better than that of “MD-AirComp (6, 20, 1), $M=1$”, which verifies the performance improvement shown in Fig. 4. Furthermore, the NMSE of “MD-AirComp (6, 20, 20), $M=4$” is the best due to more observations. Although the NMSE of “MD-AirComp (6, 20, 15), $M=8$” is smaller than -10 dB when the training rounds is larger than 200, its test accuracy is as good as the PA scheme due to the error correction property of FL.

### v2

@@ Abstract
Over-the-air computation (AirComp) is a promising technology converging communication and computation over wireless networks, which can be particularly effective in model training, inference, and more emerging edge intelligence applications. AirComp relies on uncoded transmission of individual signals, which are added naturally over the multiple access channel thanks to the superposition property of the wireless medium. Despite significantly improved communication efficiency, how to accommodate AirComp in the existing and future digital communication networks that are based on discrete modulation schemes remains a challenge. This paper proposes a massive digital AirComp (MD-AirComp) scheme that leverages an unsourced massive access protocol to enhance compatibility with both current and next-generation wireless networks. MD-AirComp utilizes vector quantization to reduce the uplink communication overhead, and employs shared quantization and modulation codebooks. At the receiver, we propose a near-optimal approximate message passing-based algorithm to compute the model aggregation results from the superposed sequences, which relies on estimating the number of devices transmitting each code sequence, rather than trying to decode the messages of individual transmitters. We apply MD-AirComp to federated edge learning (FEEL), and show that it significantly accelerates FEEL convergence compared to state-of-the-art while using the same amount of communication resources.
@@
Artificial intelligence of things (AIoT), digital over-the-air computation, unsourced massive access, federated edge learning, distributed optimization.
@@ I Introduction
With the advancement of the artificial intelligence of things (AIoT), a growing volume of data is being generated by a massive number of Internet-of-things (IoT) devices, which fuels the ongoing revolution in machine learning (ML) and artificial intelligence (AI) [2, 3, 4]. Traditionally, data collected by IoT devices is offloaded to the cloud or data centers to train massively large models [5]. However, due to concerns over data privacy and increasing computational capabilities of edge devices, AIoT in wireless networks is transitioning away from centralized learning towards distributed learning frameworks [6, 7, 5, 9, 8]. One of these emerging frameworks is federated edge learning (FEEL), where a central server coordinates multiple AIoT devices in its coverage area to train an ML model using their respective local datasets and computing resources [10]. As a result, FEEL requires solving the parallel and distributed optimization problem while considering the limited shared wireless resources and interference between devices. Model aggregation process in FEEL involves repeated uplink (UL) transmission of high-dimensional local gradients or model updates from tens to hundreds of devices, which causes a heavy burden on the multiple access networks [11].
@@ I-A Related Work
To address this problem, the concept of over-the-air computation (AirComp) has been introduced as a means of achieving communication-efficient model aggregation [10, 11, 12]. In the context of model aggregation, the objective is to compute the average of local model updates, rather than recovering each update individually. In analog AirComp, the fundamental principle involves embracing inter-user interference over the multiple-access channel (MAC) rather than mitigating it. More specifically, every element of the local model update at a device goes through pre-equalization using the inverse of the UL channel gain from the device to the BS, which acts as the parameter server. Subsequently, the pre-equalized elements are modulated onto the transmit waveform’s amplitude. Through simultaneous transmission from multiple devices, the receiver can directly obtain the sum of the local model updates from multiple devices by leveraging the superposed waveform over the MAC. To further reduce the communication overhead, by exploiting sparsification and error accumulation, the dimension of the local model updates are reduced before analog transmission in [12]. After analog AirComp, the receiver reconstructs the average of local model updates by using the approximate message passing (AMP) algorithm. A higher FEEL accuracy is achieved with limited communication resources in [13], by exploiting time-correlated sparsification on the local model updates. Multi-input multi-output (MIMO)-based AirComp is studied in [14]. Random projection technique proposed in [12] is extended to MIMO channel in [15]. Note that all the participating devices are synchronized in time in [12, 10, 11, 13, 14, 15]. If the synchronization is imperfect, [16] and [17] considered misaligned analog AirComp to solve the problem. Dynamic scheduling of devices is proposed in [18] to optimize the energy efficiency of AirComp in FEEL. We refer the readers to [19] and [20] for a comprehensive overview of FEEL and AirComp. We should emphasize that AirComp is a general technology that turns wireless medium into a joint communication and computation platform. Its benefits can also be applied to other parallel and distributed optimization applications [21]. For instance, AirComp has recently been used for efficient inference at the edge [22, 23].
@@
To fill in this gap, we propose a digital scheme that jointly enables random access and computation among massive AIoT devices. We first revisit grant-free random access (GFRA) protocols, which have been recently studied to enable massive IoT access with low signaling overhead and latency [37]. Since devices can directly transmit their signals without the permission of the base station (BS), GFRA protocols reduce the signaling overhead. There are mainly two types of GFRA schemes, i.e., sourced massive access (SMA) [39, 38, 40, 44, 42, 41, 43, 45] and unsourced massive access (UMA) [47, 46, 49, 48]. Specifically, the SMA schemes proposed in [39, 40, 41] have demonstrated their superiority in both access latency and detection performance with the help of the elegant AMP algorithm framework. Furthermore, authors in [42] and [43] extend the SMA framework to non-terrestrial networks, which efficiently improve the transmission efficiency and reduce the detection error at the same time. Both SMA and UMA schemes adopt non-orthogonal random access codebook for reducing the communication overhead. In SMA, each device has its unique non-orthogonal preamble indicating its identity. The BS performs active device detection and channel estimation, followed by data decoding [39, 38, 40, 44, 42, 41, 43, 45]. In UMA, instead, all the devices adopt the same non-orthogonal codebook. Transmitted bits are modulated by the index of the UMA codewords and decoded at the BS. The goal of the BS is to decode the list of transmitted messages, but not the identities of the active devices [47, 49, 46, 48]. We remark that the identities of active devices are not necessary for model aggregation in FEEL either. Inspired by UMA, our target is to redesign the modulation and decoding modules of conventional UMA for efficient AirComp across a large number of devices.
@@ II System Model
where $\eta$ is the global learning rate. Finally, the updated parameter vector ${\bf w}^{t+1}\in\mathbb{R}^{W}$ is sent back to the devices, and steps (2), (3), and (4) are iterated until a convergence condition is met.
@@
One reasons to have $K_{a}\ll K$ in FEEL could be due to heterogeneous computation speeds across edge devices [12, 28, 20, 53]. For a given time slot, only the devices that have finished their computation task will transmit. Alternatively devices can join or leave the learning process at different frames based on their energy or data availability as considered in typical IoT scenarios [2].
@@ III-A1 Quantization Design
We will employ a single quantization codebook to be used by all the devices, denoted by ${\bf U}\in\mathbb{R}^{Q\times N}$, where $N=2^{J}$ corresponds to $J$-bit quantization with $N$ quantization codewords, $Q\geq 1$ ($Q\in\mathbb{N}$) denotes the length of each quantization codeword, i.e., ${\bf u}_{n}\in\mathbb{R}^{Q}$, $\forall n\in[N]$. Note that $Q=1$ and $Q>1$ indicate scalar quantization and VQ, respectively.
@@ III-A2 Codebook-Based Modulation
We consider the same random access codebook shared by all the devices, denoted by ${\bf P}\in\mathbb{C}^{L\times N}$, where each column of ${\bf P}$ is a codeword of length $L$, and there are $N$ codewords in total. In the proposed scheme, there is a one-to-one mapping between ${\bf P}$ and ${\bf U}$. To be specific, if the device acquires the quantized value $[{\bf U}]_{:,n}$, then the modulation codeword (or sequence) $[{\bf P}]_{:,n}$ with the same index $n$ is transmitted. The transmission process is similar to the preamble transmission of random access procedure in 5G communication systems [24]. Hence, the proposed codebook-based modulation is more suitable for the current wireless networks compared to analog modulation typically employed in AirComp.
@@ III-A3 UL Massive Access Model
We consider a time division duplex system, where the UL and downlink (DL) channels are commonly considered the same due to channel reciprocity [40, 48]. Before each UL communication round of FEEL, the BS broadcasts a pilot signal and each active device estimates its DL channel based on this pilot signal. Then, the active devices start their grant-free UL transmissions with pre-equalization in a synchronized manner [40]. For the $d$-th element of ${\bf b}_{k}$, $\forall k,d$, the received signal ${\bf Y}_{d}\in\mathbb{C}^{L\times M}$ at the BS can be expressed as
@@
As shown in (7), the noise power can be defined by the SNR and the average power of the received signal, which is commonly used in massive access system models [37]. If the wireless channel $[{\bf h}_{k}^{\rm UL}]_{1}$ is in deep fading, the power of $[{\bf h}_{k}^{\rm UL}]_{1}$ can be extremely small. In this case, pre-equalization might increase the received signal power at the rest of the antennas. As a result, SNR at the first antenna can be too small to conduct proper signal detection. In addition, if the channel power is too small, the transmit power needed for pre-equalization might be too large. Hence, we set a threshold $\epsilon_{h}$ on the channel condition, such that, if $|[{\bf h}_{k}]_{1}|<\epsilon_{h}$, the device does not participate in this training round.
@@ IV-B Approximate Message Passing (AMP) Update Rules
where symbol $\approx$” is due to the approximation. The likelihood function and the normalization term are, respectively, denoted as
@@ IV-C2 Estimation of the activity indicators
For example, to estimate the activity indicator $a_{n,m}$, $\forall n,m$, we partially differentiate $\mathbb{E}\left\{{{\rm ln}~{}p\left({{\bf X,Y}}\right)|{\bf Y}}\right\}$ with respect to $a_{n,m}$, and let it equal zero [60]. According to (IV-B), the posterior distribution $p\left(x_{n,m}|[{\bf Y}]_{:,m}\right)$ can be approximated as $p\left(x_{n,m}|r_{n,m}\right)$, thereby simplifying the complexity of EM estimation. Hence, according to (23) and (24), the sparsity indicators $a_{n,1}$ and $a_{n,m}$, $m\in\{2,3,...,M\}$, can be obtained as follows:
@@ VI-D Results of Wireless Transmission
From Fig. 5, we can observe that NMSE decreases due to the increase in the sparsity ratio. In addition, NMSE of “MD-AirComp (6, 20, 15), $M=8$” is better than that of “MD-AirComp (6, 20, 1), $M=1$”, which verifies the performance improvement shown in Fig. 4. Furthermore, NMSE of “MD-AirComp (6, 20, 20), $M=4$” is the best due to more observations. Although NMSE of “MD-AirComp (6, 20, 15), $M=8$” is smaller than -10 dB when the training rounds is more than 200, its test accuracy is as good as the PA scheme due to the error correction property of FL.